OpenStudy (anonymous):

(logs) find the solution:10^(x+3)=6^(x)

OpenStudy (anonymous):

\[10^{x+3}=6^{x}\rightarrow x+3=x \log (6)\rightarrow x=3/(\log (6)-1)\]

OpenStudy (anonymous):

Take the common log of both sides, use the power rule on the RHS, solve for x. Not bad, eh?

OpenStudy (anonymous):

thanks